# Stabilizing Fractional Dynamical Networks Suppresses Epileptic Seizures

### v1

@@ Introduction
An accurate mathematical framework for characterizing seizure dynamics is essential for effective treatment. Dynamical network models [3] have emerged as a clear winner, allowing researchers to draw conclusions regarding the brain’s topology and function. Many works have focused on linear time-invariant dynamical networks to model and even control for epileptic activity [4, 5, 6]. However, recent evidence suggests that the brain exhibits multi-scale dynamics [7, 8], which is not captured by linear time-invariant models [9, 10, 11].
Fractional-order dynamical networks, which originated in physics and economics and quickly found their way into engineering applications [12, 13, 14, 15, 16, 17, 18], offer a middle ground between linear simplicity and multi-scale complexity. These networks accurately capture the multi-scale dynamics present in neural signals as well as the spatial relationship between brain regions [19]. Hence, fractional-order dynamical networks outperform traditional integer-order models in accurately representing neural data [20, 21, 22, 23, 24, 25, 26].
@@ Results
We assessed the goodness of model fit by computing the $R^{2}$ value for 2656 electrodes in all four epileptic brain states (interictal, pre-ictal, ictal, and post-ictal) for each patient. Segments that had an $R^{2}<0.5$ for 40% or more of their electrodes were removed from analysis to avoid unreliable results due to poor model fit (these segments were analyzed separately in the supplementary). Based on our $R^{2}$ criterion for model validation, we excluded 14 of 140 segments (10%) from analysis (see Table 1 Supplementary Material), leaving 126 good segments.
@@
After establishing an appropriate model fit, we illustrate the temporal evolution of the raw intracranial EEG (iEEG) data, goodness of fit, fractional-order exponents ($\alpha$), and eigenvalues through a complete seizure cycle of patient HUP68’s 5th seizure (Figure 2). Panel A in Figure 2 shows the iEEG recordings from 79 electrodes, with seizure onset zone (SOZ) electrodes highlighted in red. Visual inspection reveals distinct activity patterns across segments, with the ictal period showing high-amplitude, rhythmic discharges that are particularly prominent in the SOZ electrodes. In contrast, the pre-ictal period exhibits a smaller signal amplitude across all electrodes. Similarly, the SOZ electrodes during the interictal period have smaller amplitudes compared to non-SOZ electrodes.
@@
Panel C in Figure 2 shows the fractional-order exponents ($\alpha$) across all four states. Interictal fractional-order exponents are stable and tightly clustered around 0.8-1.0, while pre-ictal fractional-order exponents are smaller with a larger spread in distribution. The ictal state shows slightly larger fractional-order exponents than pre-ictal and has the most variability overall. Post-ictal fractional-order exponents are slightly larger than during ictal, with increased variability, evident in the wider IQR and outliers. Eigenvalues across all epileptic brain states are displayed in Panel D in Figure 2. Interictal eigenvalues have larger variance, with some values exceeding 1, indicating unstable or potentially marginally stable behavior. The pre-ictal eigenvalue distributions tighten, with all values less than 1, reflecting a stable system. Ictal eigenvalues have slightly larger values and a wider spread than pre-ictal, suggesting that the system is moving closer towards instability. Finally, eigenvalues during post-ictal are larger than ictal, have a greater IQR, and significantly more outliers, confirming post-seizure instability. Interestingly, the interictal state has the largest eigenvalues.
@@
Figure 9 A shows the fractional-order exponents ($\alpha$) distributions for each epileptic brain state across all patients and electrodes. Median values decreased from interictal 0.75 (IQR: 0.66-0.84) through pre-ictal 0.68 (IQR: 0.54-0.82) and ictal 0.63 (IQR: 0.48-0.77), and then increased during post-ictal 0.78 (IQR: 0.62-0.94). interictal and post-ictal segments exhibit extended tails, with interictal skewing toward smaller values and post-ictal showing spread in both directions. In contrast, fractional-order exponents ($\alpha$) during pre-ictal and ictal segments has more uniform distributions, though ictal segments show larger outliers.
@@
Several consistent patterns emerged across patients (supplementary table 2). interictal fractional-order exponents ($\alpha$) clustered around 0.70-0.80 for most patients, while pre-ictal and ictal segments remained relatively similar within individual patients. Post-ictal periods exhibited increased variability across most patients, with the exception of HUP72 and MAYO010. Six out of ten patients (HUP64, HUP68, HUP70, HUP86, MAYO011, MAYO016) showed decreased median fractional-order exponents ($\alpha$) from interictal to ictal, consistent with the group-level trajectory. Additionally, MAYO patients generally showed larger post-ictal fractional-order exponents ($\alpha$) compared to the HUP patients.
@@
Despite these overall trends, patient-level fractional-order exponents ($\alpha$) distributions varied substantially (Supplementary Figures 2 and Supplementary Table 2). Patients MAYO011 and MAYO016 showed large post-ictal values (medians = 1.02 and 0.90, respectively), whereas MAYO010 and HUP70 had consistent values across all segment types. HUP72 exhibited distinctly large values during pre-ictal and ictal segments (medians = 0.99 and 0.97).
@@
Network Stability Property Reveals Distinct Patterns Across Epileptic Brain States with High Variability during Post-Ictal
@@
Figure 9 B shows the eigenvalue distributions across all patients and electrodes. The eigenvalues showed less variability compared to fractional-order exponents ($\alpha$). Interictal eigenvalues showed the tightest distribution across all patients (median = 0.80, IQR: 0.76-0.85) while pre-ictal, ictal, and post-ictal segments exhibited similar median values (0.73, 0.72, and 0.72, respectively). However, post-ictal values showed increased variability (IQR: 0.73-0.90) compared to all other segments (pre-ictal IQR: 0.73-0.83, ictal IQR:0.72-0.80). All epileptic brain states displayed eigenvalue distributions with extended tails and numerous outliers, far greater than those of fractional-order exponents ($\alpha$). This was particularly evident in the post-ictal period. Unlike fractional-order exponents ($\alpha$), eigenvalues did not show a clear monotonic trajectory across seizure progression.
@@
Patient-level eigenvalue distributions revealed patterns consistent with group-level findings (Supplementary Figure 3, Supplementary Table 3). Eigenvalues during interictal were relatively stable across most patients (median $\approx$0.75), while pre-ictal and ictal remained similar within individual patients, mirroring the pattern observed for fractional-order exponents ($\alpha$). Eigenvalues during post-ictal periods exhibited notably increased spread, evident in the broadened violin distributions in all patients except HUP70, HUP72, and HUP78. HUP70 showed consistent eigenvalues across all brain states (consistent with its fractional-order exponents ($\alpha$) patterns), while HUP72 displayed smaller post-ictal values (median = 0.62, IQR: 0.54–0.75). Across all patients, we observe stable interictal periods, minimal variation between pre-ictal and ictal and post-ictal periods characterized by increased variability and larger values in both fractional-order exponents ($\alpha$) and eigenvalues.
@@
We performed pairwise comparisons of fractional-order exponents ($\alpha$) and eigenvalues of all epileptic brain states using population-pooled and within-patient analysis to assess the separation between states.
Population pooled analysis (figure 4) KS test showed significance between all states ($p<.001$), but with small to medium effect sizes. The largest Cohen’s $d$ values were observed for ictal vs post-ictal (0.68) and ictal vs interictal (0.65) for fractional-order exponents ($\alpha$) (panel B) and ictal vs post-ictal (0.41) and ictal vs interictal( .32) for eigenvalues (panel D). KS statistics (panels A and C) were similar for both.
Within-patient analysis (figure 5) revealed substantially larger effect sizes. For fractional-order exponents ($\alpha$) (panel B), Cohen’s $d$ value exceeded 0.9 for all comparisons except for pre-ictal vs. ictal; the largest effect sizes observed were interictal vs. ictal (d = 1.19), interictal vs. pre-ictal (d = 1.17). Eigenvalues (panel D) showed medium to large effect sizes, with interictal vs. pre-ictal ($d$ = 1.04) and interictal vs. ictal ($d$ = 1.00) showing the strongest separation. KS statistics (panels A and C) were consistently high for interictal comparisons (0.42 for all three comparisons with fractional-order exponents ($\alpha$)), indicating robust distributional separation within individual patients.
Adjacent brain states showed smaller effect sizes in both analyses. Pre-ictal vs. ictal comparisons yielded the smallest differences ($d=0.31$ for fractional-order exponents ($\alpha$), $d=0.26$ for eigenvalues), while ictal vs. post-ictal transitions showed larger effects ($d=0.95$ for fractional-order exponents ($\alpha$), $d=0.74$ for eigenvalues).
The contrast between population-pooled and within-patient effect sizes indicates substantial inter-patient heterogeneity in baseline network dynamics. While individual patients show large, consistent differences between brain states, pooling across patients reduces these effects due to varying baseline values and seizure characteristics across the cohort.
@@
Our stabilizing controller was applied at seizure onset for all ictal segments. Figure 6 illustrates the effectiveness of our control design by comparing eigenvalues in the complex plane before and after applying control in Seizure 3 of HUP68. The uncontrolled eigenvalues (red) lie near or beyond the stability boundary, while the controlled eigenvalues (blue) are well within the unit circle, satisfying the stability definition (see Methods Definition 1). Figure 7 shows the simulated controlled iEEG data compared to the original iEEG data from the seizure onset zone (SOZ) electrodes for this patient. We see that the controlled iEEG signals have substantially smaller amplitudes compared to the uncontrolled ictal iEEG signals. Similarly, figure 8 shows the percentage reduction in amplitude for all electrodes in all seizures. There was an average of $48.96\%\pm 16.94\%$ reduction in signal amplitude. There was little difference in amplitude reduction between already stable and seizures requiring control ($50.38\%\pm 18.41\%$ and $48.12\%\pm 15.95\%$, respectively).
@@ Discussion
Consistent Multi-scale and Stability Properties during interictal Periods Suggest Reproducible Brain Network Activity
The consistency in fractional-order exponents ($\alpha$) (median 0.75, IQR: 0.66-0.84) and eigenvalues (median 0.80, IQR: 0.76-0.85) across interictal periods reveals that epileptic brain networks maintain stable dynamics between seizures. These uniform distributions indicate that interictal epileptic dynamics are fundamentally non-Markovian, with neural activity substantially coupled to its temporal history [10, 23]. Although the median eigenvalue of 0.80 indicates stability, some interictal windows exhibited eigenvalues greater than 1. This may reflect model estimation uncertainty, inherent baseline instability in epileptic patients, or limitations of applying asymptotic stability to finite-time dynamics. Furthermore, within-patient analysis showed that interictal has the strongest separation from other states, with large effect sizes (Cohen’s $d>1$) for all comparisons except for interictal vs post-ictal for eigenvalues.
@@
The shift from interictal to ictal periods showed smaller values in both fractional-order exponents ($\alpha$) (0.75 to 0.63) and eigenvalues (0.80 to 0.72). A decrease in fractional-order exponents ($\alpha$) indicates stronger history-dependence of neural activity, which suggests that seizures are self-sustaining events, where the brain cannot escape from pathological activity patterns [28]. Similarly, eigenvalues also decreased in pre-ictal and ictal periods, with somewhat tightened distributions, suggesting modest increases in stability during seizures. While this appears to contradict literature that hypothesizes seizures can be mathematically represented as an instability in the brain [29, 30, 31, 32, 33], we note that asymptotic stability may not be suitable for seizures as they have finite time horizons. Notably, the stability of fractional-order systems depends on both fractional-order exponents ($\alpha$) and eigenvalues, which we show evolve together across epileptic brain states.
@@
Importantly, both fractional-order exponents ($\alpha$) and eigenvalues show small differences between pre-ictal and ictal segments. The similarity suggests that the pathological network states characteristic of seizures are already established before clinical seizure onset, consistent with previous findings that seizures do not arise instantaneously but emerge from gradual network-level transitions [34, 35]. Furthermore, the pre-ictal period may represent an already altered brain state that is primed for seizure generation rather than a simple precursor to an abrupt change [28, 36].
@@
Post-ictal fractional-order exponents ($\alpha$) and eigenvalues did not return to interictal baseline values, revealing substantial heterogeneity in recovery dynamics. Some patients exhibited values exceeding interictal baselines, suggesting hyperexcitable rebound, while others showed sustained suppression. The variability within post-ictal periods (wider IQR and numerous outliers for both metrics) indicates unstable network dynamics during recovery, where the brain may cycle through multiple states before re-establishing a baseline state. This finding indicates that post-seizure recovery is not a uniform process but rather reflects patient-specific network reorganization mechanisms [27, 37, 38].
@@
Within-patient analysis revealed substantially larger KS statistics and effect sizes compared to the population-pooled statistics. The within-patient separation between interictal and seizure-related states suggests that fractional-order parameters could serve as patient-specific biomarkers for predicting seizures, despite being limited on the population level. Furthermore, since pre-ictal and ictal states showed minimal differences, detecting the transition from interictal to pre-ictal may provide sufficient warning for impending seizures without requiring identification of seizure onset.
@@
The pattern of effect sizes reveals an asymmetry in seizure dynamics. The sharp contrast between ictal and post-ictal states ($d=1.00$ and $0.79$ for fractional-order exponents ($\alpha$) and eigenvalues, respectively) compared to the minimal pre-ictal to ictal transition ($d=0.28$ and $0.25$) indicates that seizure termination triggers more dramatic network reorganization than seizure initiation.

### v2

@@ Introduction
An accurate mathematical framework for characterizing seizure dynamics is essential for effective treatment. Dynamical network models [3] have emerged as a clear winner, allowing researchers to draw conclusions regarding the brain’s topology and function. Many works have focused on linear time-invariant dynamical networks to model and even control for epileptic activity [4, 5, 6]. However, recent evidence suggests that the brain exhibits multi-scale dynamics [7, 8], which is not captured by linear time-invariant models [9, 10, 11]. Fractional-order dynamical networks, which originated in physics and economics and quickly found their way into engineering applications [12, 13, 14, 15, 16, 17, 18], offer a middle ground between linear simplicity and multi-scale complexity. These networks accurately capture the multi-scale dynamics present in neural signals as well as the spatial relationship between brain regions [19]. Hence, fractional-order dynamical networks outperform traditional integer-order models in accurately representing neural data [20, 21, 22, 23, 24, 25, 26].
@@ Results
We assessed the goodness of model fit by computing the $R^{2}$ value for 2656 electrodes in all four epileptic brain states (interictal, pre-ictal, ictal, and post-ictal) for each patient. Segments that had an $R^{2}<0.5$ for 40% or more of their electrodes were removed from analysis to avoid unreliable results due to poor model fit (these segments were analyzed separately – see Supplementary Material). Based on our $R^{2}$ criterion for model validation, we excluded 14 of 140 segments (10%) from analysis (see Table 1 Supplementary Material), leaving 126 valid segments.
@@
After establishing an appropriate model fit, we illustrate the temporal evolution of the raw intracranial EEG (iEEG) data, goodness of fit, fractional-order exponents ($\alpha$), and eigenvalues during a complete seizure cycle of patient HUP68’s 5th seizure (Figure 2). Panel A in Figure 2 shows the iEEG recordings from 79 electrodes, with seizure onset zone (SOZ) electrodes highlighted in red. Visual inspection of iEEG data reveals distinct activity patterns across segments, with the ictal period showing high-amplitude, rhythmic discharges that are particularly prominent in the SOZ electrodes. In contrast, the iEEG data during the pre-ictal period exhibits a smaller signal amplitude across all electrodes. Similarly, the iEEG data from SOZ electrodes during the interictal period have smaller amplitudes compared to non-SOZ electrodes.
@@
Panel C in Figure 2 shows the fractional-order exponents ($\alpha$) across all four states. Interictal fractional-order exponents are stable and tightly clustered around 0.8-1.0, indicating low memory dependence and decreased multi-scale properties. Pre-ictal fractional-order exponents are smaller than fractional-order exponents during interictal and have a larger spread in distribution, indicating a higher memory dependence and increased multi-scale properties. Fractional-order exponents during the ictal period are slightly larger than the fractional-order exponents during the pre-ictal period and have the most variability overall, indicating potentially chaotic behavior and inconsistent multi-scale properties. Post-ictal fractional-order exponents are slightly larger than fractional-order exponents during ictal, with increased variability, evident in the wider IQR and outliers, indicating a shift in memory dependence and multi-scale properties.
Panel D in Figure 2 shows the eigenvalues across all epileptic brain states. Eigenvalues during interictal have larger variance, with some values exceeding 1, indicating unstable or potentially marginally stable behavior. Eigenvalue distributions during pre-ictal tighten, with all values less than 1, reflecting a stable system. Eigenvalues during ictal have slightly larger values and a wider spread than eigenvalues during pre-ictal, suggesting that the system is moving closer towards instability. Finally, eigenvalues during post-ictal are larger, have a greater IQR, and have significantly more outliers than eigenvalues during ictal, confirming post-seizure instability. Interestingly, the interictal state has the largest eigenvalues.
@@
Figure 9 A shows the fractional-order exponents ($\alpha$) distributions for each epileptic brain state across all patients and electrodes. Median fractional-order exponents decreased from interictal 0.75 (IQR: 0.66-0.84) through pre-ictal 0.68 (IQR: 0.54-0.82) and ictal 0.63 (IQR: 0.48-0.77), and then increased during post-ictal 0.78 (IQR: 0.62-0.94). Interictal and post-ictal segments exhibit fractional-order exponents with extended tails, with fractional-order exponents during skewing toward smaller values and fractional-order exponents during post-ictal showing a wider spread in both directions. In contrast, fractional-order exponents ($\alpha$) during pre-ictal and ictal segments have more uniform distributions, though fractional-order exponents during ictal segments show larger outliers.
@@
Several consistent patterns emerged across patients (see Supplementary Material Table 2). Fractional-order exponents ($\alpha$) during interictal cluster around 0.70-0.80 for most patients, while fractional-order exponents during pre-ictal and ictal segments remained relatively similar within individual patients. Fractional-order exponents during post-ictal periods exhibited increased variability across most patients, with the exception of HUP72 and MAYO010. Six out of ten patients (HUP64, HUP68, HUP70, HUP86, MAYO011, MAYO016) showed decreased median fractional-order exponents ($\alpha$) from interictal to ictal, consistent with the group-level trajectory. Additionally, MAYO patients generally showed larger post-ictal fractional-order exponents ($\alpha$) compared to the HUP patients.
@@
Despite these overall trends, patient-level fractional-order exponents ($\alpha$) distributions varied substantially (see Supplementary Material Figure 2 and Table 2). Patients MAYO011 and MAYO016 showed large fractional-order exponents during post-ictal (medians = 1.02 and 0.90, respectively), whereas MAYO010 and HUP70 had consistent fractional-order exponents across all segment types. HUP72 exhibited distinctly large fractional-order exponents during pre-ictal and ictal segments (medians = 0.99 and 0.97).
@@
Network Stability Properties Reveal Distinct Patterns Across Epileptic Brain States
@@
Figure 9 B shows the eigenvalue distributions across all patients and electrodes. The eigenvalues showed less variability compared to fractional-order exponents ($\alpha$). Eigenvalues during interictal showed the tightest distribution across all patients (median = 0.80, IQR: 0.76-0.85) while eigenvalues during pre-ictal, ictal, and post-ictal segments exhibited similar median values (0.73, 0.72, and 0.72, respectively). However, eigenvalues during post-ictal showed increased variability (IQR: 0.73-0.90) compared to all other segments (pre-ictal IQR: 0.73-0.83, ictal IQR:0.72-0.80). All epileptic brain states showed eigenvalue distributions with extended tails and numerous outliers, far greater than those of fractional-order exponents ($\alpha$), especially during post-ictal.
@@
Patient-level eigenvalue distributions revealed patterns consistent with group-level findings (see Supplementary Material Figure 3 and Supplementary Material Table 3). Eigenvalues during interictal were relatively stable across most patients (median $\approx$0.75), while eigenvalues during pre-ictal and ictal remained consistent within individual patients, mirroring the pattern observed for fractional-order exponents ($\alpha$). Eigenvalues during post-ictal periods exhibited notably increased spread, evident in the broadened violin distributions in all patients except HUP70, HUP72, and HUP78. HUP70 showed consistent eigenvalues across all brain states (consistent with its fractional-order exponents ($\alpha$) patterns), while HUP72 showed smaller eigenvalues during post-ictal (median = 0.62, IQR: 0.54–0.75). Across all patients, we observed stable eigenvalues during interictal periods, with minimal variation between eigenvalues during pre-ictal and ictal. Eigenvalues during post-ictal periods are characterized by increased variability and have larger fractional-order exponents ($\alpha$) and eigenvalues.
@@
We performed pairwise comparisons of fractional-order exponents ($\alpha$) and eigenvalues of all epileptic brain states using population-pooled and within-patient analysis to assess the separation between epileptic brain states.
Our population pooled analysis (see Figure 4) using the Kolmogorov-Smirnov statistical test showed significance between all epileptic brain states ($p<.001$), but with small to medium effect sizes. We observed the largest Cohen’s $d$ values when comparing fractional-order exponents during ictal and post-ictal (0.68) as well as during ictal and interictal (0.65) – see Figure 4 B. When comparing eigenvalues during ictal and post-ictal, we observed Cohen’s $d$ value of 0.41, and for eigenvalues during ictal and interictal, we reported a Cohen’s $d$ value of 0.32 – see Figure 4 D. Kolmogorov-Smirnov statistic values shown Figure 4 A and C were similar for both fractional-order exponents and eigenvalues.
Our within-patient analysis (see Figure 5) revealed substantially larger effect sizes compared with our population pooled analysis. For comparing fractional-order exponents ($\alpha$) shown in Figure 5 B, we observed that Cohen’s $d$ value exceeded $0.9$ for all comparisons except for fractional-order exponents during pre-ictal vs. ictal. We observed the largest effect sizes for fractional-order exponents during interictal vs. ictal ($d=1.19$) and fractional-order exponents during interictal vs. pre-ictal ($d=1.17$). When comparing eigenvalues during epileptic brain states using the Kolmogorov-Smirnov statistical test shown in Figure 5 D, we observed medium to large effect sizes. Comparing eigenvalues during interictal vs. pre-ictal showed Cohen’s $d=1.04$, and comparing eigenvalues during interictal vs. ictal showed Cohen’s $d=1.00$, which showed the strongest separation (see Figure 5 D). Kolmogorov-Smirnov statistics shown in Figure 5 A showed consistently large Cohen’s $d$ values for fractional-order exponents during interictal in comparison with all other epileptic brain states (0.42 for all three comparisons), indicating robust distributional separation within individual patients. Kolmogorov-Smirnov statistics shown in Figure 5 D showed consistent Cohen’s $d$ values for eigenvalues as observed in Figure 5 A for fractional-order exponents.
Adjacent epileptic brain states (i.e., interictal vs. pre-ictal, pre-ictal vs. ictal, and ictal vs. post-ictal) showed smaller effect sizes in both individual and population analyses. Both fractional-order exponents and eigenvalue comparisons during pre-ictal vs. ictal yielded the smallest effect size with $d=0.31$ for fractional-order exponents and $d=0.26$ for eigenvalues. On the other hand, both fractional-order exponents and eigenvalue comparisons during ictal vs. post-ictal showed larger effect sizes with $d=0.95$ for fractional-order exponents and $d=0.74$ for eigenvalues.
The contrast between population-pooled and within-patient effect sizes suggests substantial inter-patient heterogeneity in baseline epileptic dynamics. While individual patients show large, consistent differences between epileptic brain states, pooling across patients reduces these effects due to varying baseline values and seizure characteristics across the cohort.
@@
Our stabilizing controller was applied at seizure onset for all ictal segments. Figure 6 illustrates the effectiveness of our control strategy by comparing eigenvalues in the complex plane before and after applying control in Seizure 3 of HUP68. The uncontrolled eigenvalues (red) lie near or beyond the stability boundary, while the controlled eigenvalues (blue) are well within the unit circle, satisfying the stability definition (see Methods Definition 1). Figure 7 shows the simulated controlled iEEG data compared to the original iEEG data from the seizure onset zone (SOZ) electrodes for Seizure 3 of HUP68. We see that the controlled iEEG signals have substantially smaller amplitudes compared to the uncontrolled ictal iEEG signals. Figure 8 shows the percent reduction in amplitude for all electrodes in all 35 seizures. We observed an average reduction of $48.96\%\pm 16.94\%$ in signal amplitude. There was little difference in amplitude reduction between fractional-order systems that were already stable during ictal and unstable fractional-order systems requiring control ($50.38\%\pm 18.41\%$ and $48.12\%\pm 15.95\%$, respectively).
@@ Discussion
Consistent Multi-scale and Stability Properties Across Patients Suggest Reproducible Brain Network Activity
Our analysis of fractional-order exponents and eigenvalues across seizure states reveals that epileptic brain networks exhibit consistent, multi-scale memory dynamics and structured stability properties. Fractional-order exponents quantify history-dependence in neural dynamics, a property that integer-order models cannot capture. These fractional-order exponents reveal how strongly current brain states depend on past activity across multiple timescales, with lower values indicating stronger memory dependence and higher values indicating weaker memory dependence, and values approaching 1 reducing to the integer-order case. Across patients, we observed a reproducible pattern across epileptic brain states. Interictal fractional-order exponents indicate low memory dependence and a decrease in multi-scale properties. Pre-ictal fractional-order exponents reflect a higher memory dependence and an increase in multi-scale properties. Fractional-order exponents during the ictal period indicate potentially chaotic behavior and inconsistent multi-scale properties. On the other hand, post-ictal fractional-order exponents indicate a shift in memory dependence and multi-scale properties, without returning to interictal levels. Similarly, eigenvalues during interictal indicate unstable or potentially marginally stable behavior. Pre-ictal eigenvalue distributions generally reflect stabilized behavior, while the ictal eigenvalues suggest that the brain is moving closer towards instability. Finally, eigenvalues during post-ictal indicate instability, suggesting disrupted or incomplete recovery following seizure termination.
In Figure 9 A, the interictal segments exhibit the weakest memory dependence (highest $\alpha$), while pre-ictal and ictal periods show strong memory effects. Post-ictal fractional-order exponents indicate that these segments are much more chaotic and do not return to interictal levels. This trend is mirrored in Figure 9 B with the eigenvalues and stability. The consistency in fractional-order exponents ($\alpha$) (median 0.75, IQR: 0.66-0.84) and eigenvalues (median 0.80, IQR: 0.76-0.85) across interictal periods reveals that epileptic brain networks maintain consistent dynamics between seizures. These uniform fractional-order exponent distributions indicate that interictal epileptic dynamics are fundamentally non-Markovian, with neural activity substantially coupled to its temporal history [10, 23]. We consistently observed eigenvalues greater than 1 during interictal periods, indicating unstable dynamics. Our within-patient analysis revealed that eigenvalues during interictal have the strongest separation from other epileptic brain states, with large effect sizes (Cohen’s $d>1$) for all comparisons except for eigenvalues during interictal vs post-ictal.
These patterns and strong effect sizes highlight the structured nature of seizure dynamics, aligning with prior studies that have used EEG data to track physiological changes [28]. Our findings are also consistent with the theory of critical slowing, which has been proposed as a biomarker for seizure susceptibility [29].
Seizure onset and termination have been modeled as critical transitions in dynamical systems [30, 31, 29, 32, 33], and recent work has explored seizure dynamics across spatial scales [30, 34]. Other studies have identified self-similar properties in EEG seizure data using multivariate eigen-wavelet analysis [35], generalized Hurst exponents to distinguish healthy from epileptic activity [36], and fractal features for machine learning classification in a clinical context [37]. Our work echoes the conclusions of these previous studies by explicitly modeling scale-dependent memory and stability, offering a unified framework to characterize and track multiscale seizure dynamics across time.
@@
The shift from interictal to ictal periods showed a decrease in fractional-order exponents ($\alpha$) (0.75 to 0.63) and eigenvalues (0.80 to 0.72). Lower fractional-order exponents ($\alpha$) indicates stronger history-dependence of neural activity, which suggests that seizures are self-sustaining events, where the brain cannot escape from pathological activity patterns [38]. Similarly, eigenvalues also decreased in pre-ictal and ictal periods, with somewhat tightened distributions, suggesting a modest shift toward a more robust, stable region during seizures. While this observation may initially appear to contradict recent literature that hypothesizes that seizures can be mathematically represented as an instability in the brain [39, 40, 33, 41, 42], we note that asymptotic stability (assessed in this work) may not be suitable for characterizing seizure onset since seizure events have finite time horizons. Notably, the stability of a fractional-order dynamical network depends on both its fractional-order exponents ($\alpha$) and its spatial matrix (AA). In this work, we observe that fractional-order exponents evolve together with eigenvalues across epileptic brain states.
@@
Importantly, both fractional-order exponents ($\alpha$) and eigenvalues show small differences between pre-ictal and ictal segments. The similarity between pre-ictal and ictal suggests that the pathological network states characteristic of seizures are already established before clinical seizure onset, consistent with previous findings that seizures do not arise instantaneously but emerge from gradual network-level transitions [43, 44]. Furthermore, the pre-ictal period may represent an already altered brain state that is primed for seizure generation rather than a simple precursor to an abrupt change [38, 45].
@@
Post-ictal fractional-order exponents ($\alpha$) and eigenvalues did not return to interictal baseline values, revealing substantial heterogeneity in recovery dynamics. Some patients exhibited both fractional-order exponents and eigenvalues exceeding interictal baselines, suggesting hyperexcitable rebound, while others showed sustained suppression. The variability within post-ictal periods (wider IQR and numerous outliers for both metrics) indicates unstable network dynamics during recovery, where the brain may cycle through multiple states before re-establishing a baseline state. This finding indicates that post-seizure recovery is not a uniform process but rather reflects patient-specific network reorganization mechanisms [27, 46, 47].
@@
Within-patient analysis revealed substantially larger Kolmogorov-Smirnov test statistics and effect sizes compared to the population-pooled statistics. The strong within-patient separation between interictal and seizure-related states suggests that fractional-order exponents could serve as patient-specific biomarkers for predicting seizures, despite being limited on the population level. Additionally, since fractional-order exponents during pre-ictal and ictal states showed minimal differences, detecting the transition from interictal to pre-ictal may provide sufficient warning for impending seizures without requiring the exact time of seizure onset. This is further supported by the observation that interictal to ictal, and interictal to pre-ictal comparisons exhibited similarly strong separability, as shown in Figure 5.
@@
The pattern of effect sizes reveals an asymmetry in seizure dynamics. The sharp contrast between ictal and post-ictal states ($d=1.00$ for fractional-order exponents and $0.79$ for eigenvalues) compared to the minimal difference between pre-ictal and ictal states ($d=0.28$ for fractional-order exponents and $0.25$ for eigenvalues) may indicate that seizure termination triggers more dramatic network reorganization than seizure initiation. This finding aligns with recent modeling work showing that seizure onset and spread can emerge as critical transitions in brain networks driven by changes in excitability and connectivity [32]. Furthermore, our finding that fractional-order exponents and eigenvalues separate interictal and pre-ictal states is supported by prior work highlighting the importance of identifying network-level changes as patient-specific biomarkers for seizure prediction [48].
@@
To our knowledge, our work represents one of the most comprehensive demonstrations of fractional-order control for suppressing seizure dynamics applied to real patient iEEG data. Our stabilizing control framework achieved seizure suppression in 34/35 seizures, successfully stabilizing 77% of initially unstable seizures and reducing seizure amplitude by approximately 49% across all electrodes. Notably, our control strategy produced a similar average amplitude reduction in both initially stable and unstable seizures, showcasing effective suppression of ictal activity regardless of mathematical stability status. These findings suggest that fractional-order control strategies can generalize across diverse patient pools.
In contrast, previous work has focused on developing state-feedback stabilization strategies for linear time-invariant models of physiological networks [49]. Another example is a generalized pole placement algorithm for linear time-invariant dynamics [4]. The main issue with linear-time invariant models is that they do not capture the measurement dependence inherent in physiological networks [19].
Recent work has focused on designing linear networked models [50, 51, 52]. For example, the work in [50] derives graph-theoretic conditions for structural vibrational stabilizability under which linear networked models can always be stabilized, and the work in [51] provides a method to design the vibrational inputs to stabilize the linear network. The work in [52] studies interconnected excitatory-inhibitory pairs with linear threshold dynamics, and presents strategies to design networks with desired robustness properties. These methods may offer theoretical guarantees under ideal conditions, but are not easily estimated from data and may not accommodate the heterogeneity observed across patients.
Open-loop vibrational control strategies have been proposed to stabilize networks of nonlinear oscillators [53, 54, 55]. For example, recent work shows that vibrational control synchronizes clusters of nonlinear oscillators in a network by providing sufficient conditions for uniform exponential stability of a fixed point  [53]. Furthermore, the work in [54] focuses on stabilizing a nonlinear network of oscillators by providing the patterns of the pairwise relationships between the oscillators’ phases. In addition, the work in [55] provides sufficient conditions for vibrational inputs to stabilize cluster synchronization of a nonlinear network of oscillators and offers a tractable approach for designing vibrational control. Finally, for a bistable dynamical system, the authors provide conditions on external perturbations to ensure input-to-state stability [40]. The main limitations of considering a network of nonlinear oscillators or bistable dynamical systems is the inability to estimate the parameters of the model from data and the issue of scalability.
Fractional-order dynamical networks can be efficiently estimated from data [56, 57, 58] and do not suffer from issues of scalability. In our prior work, we derived conditions for global asymptotic stability of time-invariant fractional-order systems in discrete-time [59, 60]. Furthermore, we derived two stabilizing feedback strategies by altering the parameters of fractional-order systems and provided the polynomial-time computational complexity to compute these feedback strategies [60].
In this work, we demonstrate the capability of our state-of-the-art stabilizing state feedback control scheme to effectively suppress epileptic activity in 34 out of 35 patients. Future work will focus on comparing our state feedback control method with traditional stabilizing state feedback methods for linear time-invariant dynamics.